OpenStudy (anonymous):

Simplify- will medal and fan (−2⋅qp^3)^3 Select one: a. m^3⋅p^12⋅q^6 b. 27⋅q^3⋅p^9 c. 4⋅m^8⋅p^6⋅q^8 d. −8⋅q^3⋅p^9

OpenStudy (campbell_st):

the external power of 3 operates on each part inside the brackets \[(-2)^3 \times (p)^3 \times (q^3)^3\]

OpenStudy (campbell_st):

you should be able to simplify this

OpenStudy (anonymous):

@AWillie98 do you sill need help?

OpenStudy (anonymous):

yes @Miss.SweetiePie

OpenStudy (anonymous):

\[(-2)^{3}= -2 \times -2 \times -2 = -8\]

OpenStudy (anonymous):

\[(p)^{3}=p ^{3}\]

OpenStudy (anonymous):

\[(q ^{3})^{3}=q ^{9}\]

OpenStudy (anonymous):

which gives you D

OpenStudy (anonymous):

thank you @Miss.SweetiePie

OpenStudy (anonymous):

any time

OpenStudy (campbell_st):

nice thing about this site, if you wait long enough someone always give the answer

OpenStudy (anonymous):

@campbell_st who are you talking about????

OpenStudy (campbell_st):

the mantra of open study.... help improve understanding and don't give answers

OpenStudy (campbell_st):

@annas, perhaps the question was to difficult for open study

OpenStudy (campbell_st):

I choose to help for free... and many do... that way I can pick and choose... @annas